Suppose we have a production function for wheat 

 Q = 𝐾𝐿 − 0.8K2 − 0.2𝐿

2

A. Given K = 10 Units, calculate the level of labour that the AP reaches max and how 

much wheat is produced at that point

B. Given K = 10 at what level of labour output reaches maximum

C. How much level of labor should the company hire in order to be efficient


1
Expert's answer
2021-05-25T11:14:55-0400

Given,

Production function:

"Q=KL\u22120.8K^\n\n2\n\n\u22120.2L\n\n^2"

(A).

"Q=10L\u22120.8(10)^\n\n2\n\n\u22120.2L^\n\n2"

"Q=10L\u22120.8\u00d7100\u22120.2L\n\n^2"

"Q=10L\u221280\u22120.2L^\n\n2"

"AP=\frac{Q

}{L}"


"AP=\frac{10L−80−0.2L^

2}{

L}"


"AP=10−\frac{80
}{
L}



−0.2L"

AP will be maximum if the first order derivative is zero

"\frac{dAP

}{dL}



=−(\frac{−80}{

L

^2



})−0.2"


"\frac{dAP}{

 dL}



=\frac{80}{

L

^2}



−0.2"


"\frac{dAP

}{dL}



=0"


"\frac{80}

{L^

2
}


−0.2=0"


"\frac{80
}{
L^

2



}=0.2"


"L^

2

=\frac{80}{

0.2

}"

"L^\n\n2\n\n=400"

"L=20"


L=20 will give maximum AP.


Substitute K=10 in the production function.


The quantity of wheat at L=20 is given below:

"Q=10\u00d720\u221280\u22120.2\u00d7(20)^\n\n2"

"Q=200\u221280\u221280"

"Q=40"

40 units of wheat will be produced.


(B).

To find maximum output we will find the first-order derivative (slope) of the production function and set that equal to zero.

"\frac{dQ

}{dL}



=\frac{d(10L−80−0.2L^

2

)

}{dL}"


"\frac{dQ

}{dL



}=10−0.2×2L"


"\frac{dQ

}{dL



}=10−0.4L"


"\frac{dQ

}{dL



}=0"


"10\u22120.4L=0"


"10=0.4L"


"L=\frac{10

}{0.4}"


"L=25"

At L=25 the output will be maximum.


(c).

Since L=25 gives the maximum output the firm will hire 25 units of labor to be efficient. After L=25 the total production will decrease and it will be inefficient for the firm to produce above L=25.